OpenStudy (anonymous):

Please help . I have ten minutes left to work on these! What is the GCF of 24a^3b and 36ab^2? 2ab 4a^3b 12ab 72a^3b^2 Thank You so much!! :)

OpenStudy (raden):

24 = 2^3 * 3 36 = 2^2 * 3^2 so, the GCF of 24 and 36 = 2^3 * 3^2 = ... then how to get the GCF a^3b and ab^2 ? just take the highest degrees, it is a^3b^2 then multiply both GCF above

OpenStudy (anonymous):

literally, i have ten minutes to end semester. please give answer. i have 4 problems left!

OpenStudy (raden):

2^3 * 3^2 = 8 * 9 = 72 so, it is 72a^3b^2